OpenStudy (anonymous):

Solve the system by substitution -4.5x-2y=-12.5 3.25x-y=-0.75 Please help me! I suck at substitution. Can you show me step by step to help me understand?

OpenStudy (gorv):

3.25x-y=-0.75 adding 0.75 both side 3.25x-y+0.75=0 adding y both side 3.25x+0.75=y put this value of y in other equation

OpenStudy (gorv):

-4.5x-2y=-12.5 -4.5x-2*(3.25x+0.75)=-12.5 now solve it and tellme what u got